Complete working instructions found
PRIME FOCUS ROTATOR:
Steve Magee has increased the motor torque on the INT prime focus rotator
and after many trails it appears to move reliably between the 0 and
180 deg position. There is not provision made in the microcomputer
software to indicate the TCS that the locking pin (PLUNGER) is in and
this has caused problems in the past. When using the PF rotator I strongly
suggest you use the following procedure.
- ALWAYS OPERATE THE ROTATOR FROM THE TCS CONSUL type either USER> rot
sky 0 or USER> rot sky 180 and wait 1 minute for the action to complete.
- In the green "Prime Rotator Tip Session" window (lpss13 - workspace
/program window) type PLUNGER_IN after a
couple of seconds you should get a message SUCCESS CODE = N it is then
safe to continue observing. If SUCCESS CODE = T (time out) then execute
- In the green "Prime Rotator Tip Session" window type PLUNGER_OUT
followed by CLAMP_OFF assuming you have a SUCCESS CODE = N for both of
these actions then type PLUNGER_IN and after a SUCCESS CODE = N type
CLAMP_ON and it should then be safe to continue observing.
Eng. Notes for moving the rotator:
-Telescope in acces park.
-Plug the VT200 in the PF connector, "Telescope Top Ring"
VT200 Protocol: 7 Bits, 9600 Baud, 7 Bits odd Parity no check, 1 Stop Bit,
No Local Echo.
-Type HELP, Will display Instructions, Before activating a mechanism you must
select it by typing ROTATION
-Before moving the rotation you must free the rotator.
TO MOVE ROTATOR
XXXX POSITION where XXXX is the number of a position.
0 = 200
90 = 1100
180 = 2000
270 = 2900
360 = 3800
Position may be obtained by typing:
?CARDINAL informs if at cardinal point.
N - No error detected
O - Mechanism is operating and its status may not be meaningfull
C - Command error invalid mechanisn identifier has been set.
T - Timeout error, Taking longer than it should to reach posn.
L - Limit error reach a limit switch. Should not occur in normal operation.
S - Switch error, both limit switches are true ie. cable disconnected.
D - Demand error, mechanism has been commanded to move outside its normal
-The electronics 4ms controller and the Power Supply is in the clip center.
Loading the RAM disk into the WFC controller:
The RAM disk of the WFC controller may be corrupted if neither
rebooting (tip meng) or power cycling gets the connection back
properly. (Check with ccd0>see, 1 tele, 2 tele, 3 tele 4 tele).
In that case, the firmware of the eprom is corrupted,
and you'll have to read it back from the floppy disk.
IMPORTANT: Before powercycling the controller, switch off the
portserver located below the blue box on the cass rotator. Switch
back on after the controller has been powercycled.
Remove the floppy drive, check with 'see' if everything is ok
- Connect the gantry terminal to the VDU socket on the ccd controller
- Find portable floppy drive + correct floppy
- Switch off controller
- Plug blue ribbon calbe in back of controller
- Switch on floppy drive + controller
- On gantry terminal, press CR to get ok> prompt
- ok> insite 1 drive
- ok> 100 restore-eeprom (100 is block number, see disk)
- ok> reboot
It is also possible to reload the RAM from the tip meng session
NB:Do not attempt this unless you
are absolutely sure of what you are doing since loading the wrong
software could destroy the science CCDs.
There are a number of defective pixels on the INT Autoguider CCD. It is
easy to confuse these with stars as they will be picked up by a FIELD
command. The main difference though is that they don't move.
Below is a list of the pixel coordinates of all the defects I'm aware of:-
X Y Mag
--------- ------- ---------
17 746 16
17 911 17
17 981 16
17 1000 16
110 800 13
237 680 16
275 702 16
326 776 18
414 731 17
420 897 17
564 929 16
637 930 16